A typical city water tower holds on average 80000L of water. This equates to a mass of 8.0 x 107 grams. What percent of this mass comes from oxygen?


1
Expert's answer
2020-02-17T07:06:39-0500

First, the number of moles of water in a given volume should be calculated:

n = m/Mr

where n - number of moles, m - mass, Mr - molecular weight.

From here,

n (water) = 8.0 × 107 g / 18 g/mol = 4.4 × 106 mol

Second, as one water molecule contains one oxygen atom, the number of moles of oxygen in the given volume is the same. As a result, the mass of oxygen equals:

m (oxygen) = n × Mr = 4.4 × 106 mol × 16 g/mol = 7.1 × 107 g

Finally, the percentage of oxygen in a given mass of water is:

w = m (oxygen) / m (water) = 7.1 × 107 g / 8.0 × 107 g = 0.88875 = 89%


Answer: 89%